# Imperative Learning: A Self-supervised Neuro-Symbolic Learning Framework for Robot Autonomy

### v5

@@ 5.4.6 Effectiveness Validation
Table 15: The percentage of ATE decreasement of the VO network in iSLAM after self-supervised training on half sequences in KITTI and testing on the other half.
Trained on Tested on Enhancement
01, 02, 04, 05, 06 00, 07, 08, 09, 10 14.0%
00, 01, 04, 06, 08 02, 05, 07, 09, 10 42.9%
00, 06, 08, 09, 10 01, 02, 04, 05, 07 25.7%
00, 02, 08, 09, 10 01, 04, 05, 06, 07 43.4%
5.4.7 Adaptation to New Environments
Since iSLAM is a self-supervised learning framework, we can achieve online learning: as no labels are required, the network can learn to adapt to new environments while performing tasks in it. To validate this hypothesis, we conduct an experiment where the VO model is trained on several random halves of the sequences in the KITTI dataset and subsequently tested on the remaining halves. The results are shown in Table 15. It is seen that the ATE reduced 14%-43% after self-supervised training when compared to the pre-train model. Note that the pre-train model has never been trained on KITTI before. This suggests that imperative learning enables the VO network to acclimate to new environments by allowing it to acquire geometry knowledge from the back-end.
@@ 5.5.1 Baselines
(a) RL baseline on instance #1.
(b) OR-Tools with $1800\mathrm{s}$ runtime on instance #1.
(c) iMTSP on instance #1.
(d) RL baseline on instance #2.
(e) OR-Tools with $1800\mathrm{s}$ runtime on instance #2.
(f) iMTSP on instance #2.
Figure 17: The figures visualize the performance of the baselines and iMTSP on two MTSP instances with $5$ agents and $500$ cities, where instance #1 has a central depot and instance #2 has an off-center depot. The numbers denote the length of the route, and different colors represent different routes. The longest tour lengths are underlined in each figure. iMTSP produces the best solutions and has fewer sub-optimal patterns like circular partial routes or long straight partial routes.
@@ 5.5.6 Qualitative Performance
We next conduct qualitative analysis for the MTSP solvers and show two representative instances in Figure 17, where different colors indicate different routes. It is observed that sub-optimal patterns exist for both baselines but very few of them exist in that of iMTSP. For example, we can observe circular partial routes for Google OR-Tools such as the green and purple routes of instance #$\#1$, where the route length can be reduced by simply de-circle at the overlapping point. Similarly, long straight partial routes are present in blue and purple in the instance #$\#2$ of the RL baseline, which indicates those agents travel for a long distance but only visit $2$ cities, e.g., the beginning and the terminal city of the partial route. These observations provide further intuitions about the advantages of iMTSP.
@@ 6 Conclusions & Discussions
We believe that the theoretical challenges of imperative learning will also inspire new directions and topics for fundamental research in bilevel optimization. For example, second-order bilevel optimization remains underexplored in machine learning due to time and memory constraints. However, its relevance is increasingly critical in robotics, where second-order optimization is essential for achieving the required accuracy in complex tasks. Furthermore, handling lower-level constraints within general bilevel optimization settings presents significant challenges and remains an underdeveloped area. Recent advancements, such as those based on proximal Lagrangian value functions (Yao et al., 2024, 2023), offer potential solutions for constrained robot learning problems. We intend to investigate robust approaches and apply these techniques to enhance our experimental outcomes for imperative learning. Additionally, we plan to develop heuristic yet practical solutions for bilevel optimization involving discrete variables, and we aim to refine the theoretical framework for these discrete scenarios with fewer assumptions, leveraging recent progress in control variate estimation.
@@
This work was, in part, funded by DARPA grant DARPA-PS-23-13. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of DARPA.

### v6

@@ 5.5.1 Baselines
(a) RL baseline.
(b) OR-Tools with $1800\mathrm{s}$ runtime.
(c) iMTSP.
Figure 17: The performance visualization of the baselines and iMTSP for a problem instance with a central depot, $5$ agents, and $500$ cities. The numbers denote the length of the route, and different colors represent different routes. The longest tour lengths are underlined and iMTSP has the best solution and has fewer sub-optimal patterns like circular partial routes or long straight routes.
@@ 5.5.6 Qualitative Performance
We next conduct qualitative analysis for the MTSP solvers in Figure 17, where different colors indicate different routes. It is observed that sub-optimal patterns exist for both baselines but very few of them exist in that of iMTSP. For example, we can observe circular partial routes for Google OR-Tools such as the green and purple routes, where the route length can be reduced by simply de-circle at the overlapping point. This observation provides further intuitions about the advantages of iMTSP.
@@ 6 Conclusions & Discussions
We believe that the theoretical challenges of imperative learning will also inspire new directions and topics for fundamental research in bilevel optimization. For example, second-order bilevel optimization remains underexplored in machine learning due to time and memory constraints. However, its relevance is increasingly critical in robotics, where second-order optimization is essential for achieving the required accuracy in complex tasks. Furthermore, handling lower-level constraints within general bilevel optimization settings presents significant challenges and remains an underdeveloped area. Recent advancements, such as those based on proximal Lagrangian value functions (Yao et al., 2024a, 2023), offer potential solutions for constrained robot learning problems. We intend to investigate robust approaches and apply these techniques to enhance our experimental outcomes for imperative learning. Additionally, we plan to develop heuristic yet practical solutions for bilevel optimization involving discrete variables, and we aim to refine the theoretical framework for these discrete scenarios with fewer assumptions, leveraging recent progress in control variate estimation.
@@
This work was supported by the DARPA award HR00112490426. Any opinions, findings, conclusions, or recommendations expressed in this paper are those of the authors and do not necessarily reflect the views of DARPA.